# Tidal heating in binary inspiral of strange quark stars

### v1

@@ I Introduction
All these previous works considered only the hadronic sources of shear viscosity originating from $ee$ scattering as the dominant source of shear viscous damping at low temperatures ($T\ll 10^{10}$ K which is $\approx 1$ MeV) during the inspiral. But, as explained earlier, the densities at the core of neutron stars can be high enough for strange matter such as hyperons or deconfined quark matter to appear as stable components. Recently, Ghosh et al. Ghosh et al. (2024) demonstrated that the bulk viscosity originating from non-leptonic weak processes involving hyperons can be several orders higher ($\approx 10^{8}-10^{10}$ times) than shear viscosity from $ee$ scattering in the temperature range of $10^{6}-10^{8}$indblom and Owen (2002); Chatterjee and Bandyopadhyay (2006); Alford and Haber (2021), and the dissipation of the dominant flimit-$f-$mode introduces an additional phase of the order $0.510^{-3}-$ 0.5 rad to the GW signal depending on the neutron star masses, which should in principle be detectable by the next generation GW detectors (Ghosh et al., 2025).
@@
In general, viscous dissipation in strange quark stars is driven by shear and bulk viscosities. Shear viscosity generated by quark-quark scattering due to gluon exchange has been calculated up to $\mathcal{O}(\alpha_{s})$ in a standard Fermi liquid approximation Haensel and Jerzak (1989) and also for degenerate quark matter where dynamical screening unfolds Heiselberg and Pethick (1993). At typical temperatures and densities in compact stars, shear viscosity results to be orders of magnitude smaller than the bulk viscosity, the latter being more relevant in the tidal energy dissipation.
@@
This article is structured as follows: in Sec. II we briefly recapitulate the Newtonian tidal heating formalism during the inspiral stage of the coalescence of BNS and estimate the bulk viscous dissipated energy. We also present the non-ideal bag model we use to describe unpaired strange quark matter in subsection II.1, the strange star EoS in subsection II.2, and calculate the bulk viscosity associated with non-leptonic weak interactions among quarks in subsection II.3, which are used to calculate the tidal dissipated energy in binary inspiral. In Sec. III, we estimate the temperature reached during the inspiral due to this dissipation from strange quark bulk viscosity. Then in Sec.IV, we estimate the additional phase correction introduced in the frequency domain gravitational waveform due to this viscous dissipation and compare with the current limits with future generation GW detector sensitivity to argue about the detectability. In Sec. V we discuss the main implications of this work and also future directions. We use natural units, assuming $G=c=1$, unless explicitly stated otherwise.
@@ II Tidal dissipated energy from mode damping
$\gamma_{\alpha}=\dot{E}_{\text{visc},\alpha}/2E_{\alpha},$ (1)
@@
where $\xi_{\alpha}(r)$ is the eigenfunction and $a_{\alpha}(t)$ is the amplitude of the particular eigenmode due to the tidal field of the companion and as stated before, $\alpha$ is the normal mode index. For a particular mode, the eigenfunction can be written as the sum of the radial $\xi^{r}_{nl}$ and tangential $\xi^{\perp}_{nl}$ components
@@
[…]
where $\xi^{r}$ and $\xi^{\perp}$ are the radial and perpendicular components of the $f$-mode eigenfunction, and the corresponding velocity field can be written as $\bm{v}=-i\omega\bm{\xi}$, where $\omega$ is the $f$-mode angular frequency, which is determined via a relativistic Cowling approximation Pradhan and Chatterjee (2021) as outlined in appendix A. The normalised mode eigenfunctions are also used to calculate the tidal coupling $Q_{nl}$ defined as
[…]
where $R$ is the radius of the star and $\rho$ is the energy-density. In the adiabatic limit away from the merger when the orbital evolution is sufficiently slow ($\dot{\Omega}/\Omega\ll 1$ where $\Omega$ is the orbital angular velocity), and far away from the merger and the resonant flimit-$f-$mode frequency, we can estimate the energy dissipated as Ghosh et al. (2024)
$\dot{E}_{\rm visc}=\frac{24\pi}{5}q^{2}(1+q)\frac{M_{1}^{2}}{R}\frac{Q^{2}}{\bar{\omega}^{4}}\left(\frac{R}{D}\right)^{9}\gamma,$ (9)
where $M_{1}$ and $qM_{1}$ are the mass of the heavier star and its companion respectively with $q$ being the mass ratio ($q\leq 1$), $Q$ tidal coupling strength of the $l=m=2$ $f$-mode, $R$ is the radius of the star, $\bar{\omega}$ the normalised (by $\sqrt{M/R^{3}}$) angular frequency of the $f$-mode and $D$ is the orbital separation.
II.1 EoS for strange quark matter
We briefly describe the non-ideal bag model that we use to study unpaired strange quark matter ($u,d,s$) with electrons $e$. The thermodynamic potential $\Omega_{\rm ni}$ (normalised by the volume) as a function of the chemical potential of the constituent particles $\mu_{i}$ with $i=u,d,s,e$ and temperature $T$ can be written as follows
$\Omega_{\rm ni}=\Omega_{e}^{(0)}+a_{4}\sum_{f=u,d,s}\Omega_{i}^{(0)}+B_{\rm eff},$ (10)
@@ II.1 EoS for strange quark matter
where $E_{i,k}=\sqrt{k^{2}+m_{i}^{2}}$ and $g_{i}$ is the degeneracy factor and $m_{i}$ is the mass of the particles. For electrons, $g_{e}=2$, considers their spin and for quarks, $g_{f}=2N_{c}$, includes the spin and color charge degrees of freedom being $N_{c}=3$ the number of colors. The mass of the electron is around $0.511$ MeV, but here we consider it to be massless. In this approximation, the expression in Eq. (11) is given by
@@
$+$ $\left.\frac{3}{2}m_{f}^{4}\ln{\left(\frac{\mu_{f}+k_{f}}{m_{f}}\right)}\right],$
@@
and the beta equilibrium conditions at low temperature
@@
Figure 1: Stability window for the non-ideal bag model. The strange quark mass $m_{s}$ is plotted as a function of $B_{\rm eff}^{1/4}$ for three different values of the nonperturbative parameter $a_{4}$. The shadow regions represent stable strange quark matter according to the Bodmer-Witten conjecture. We have imposed $m_{s}=95$ MeV as the lowest value of the strange quark mass.
II.2 Strange Star EoS and Stellar Properties
@@
In Fig. 1, we plot the regions which satisfy the constraints imposed by the Bodmer-Witten conjecture in the parameter space of $B_{\rm eff}^{1/4}$ and $m_{s}$ as reported in Ref. Torres and Menezes (2013). The vertical lines represent the region where the energy density per baryon for two-flavour quark matter equals $930$ MeV and the other curved lines depict the values where the energy density per baryon for strange quark matter is $930$ MeV. The areas to the left of the vertical lines represent the regions where two-flavour quark matter is stable for a given value of $a_{4}$. The shaded regions show the stability windows setting $m_{s,\text{min}}=95$ MeV as the lowest value for the strange-quark mass and $m_{u}=m_{d}=4$ MeV.
[…]
According to the stability windows found in Fig. 1, we define the sets of parameters ($m_{s}$, $a_{4}$ and $B_{\rm eff}^{1/4}$) in table 1 so as to determine the EoSs, while assuming $m_{u}=m_{d}=4$ MeV and massless electrons. Then, we solve the Tolman-Oppenheimer-Volkoff equations (TOV equations) to obtain the mass (M)- radius (R) relations and the $f$-mode oscillation frequencies within the Cowling approximation as described in the appendix A. In Fig. 2, we display the M-R relations obtained with the set of parameters described above. Comparing the curves obtained using sets A-B, C-D, and E-F, where $a_{4}$ and the bag constant are fixed, we infer that decreasing the value of the strange quark mass produces higher maximum masses and radii. Using the EoS of the non-ideal bag model it is possible to produce maximum masses above $2M_{\odot}$ unlike the original MIT bag model, which generates at most a maximum mass up to $1.85M_{\odot}$ Lopes et al. (2021). A similar effect that is more remarkable occurs when we decrease $a_{4}$ Lopes et al. (2021) (see, for example, set A and C, the small variation of the bag constant does not have a significant effect).
@@
Figure 2: M-R sequences for the EoS parametrization tabulated in table 1. Horizontal bands correspond to masses $M=2.08{\pm 0.07}M_{\odot}$ of PSR J0740$+$6620 (Fonseca et al., 2021) and $M=2.01^{+0.04}_{-0.04}M_{\odot}$ of PSR J0348$+$0432 (Antoniadis et al., 2013). The 90% contour of M-R measurement for PSR J0740+6620 corresponding to Riley et al. (Riley et al., 2021) is shown in black, and for PSR J0030+0451 (Riley et al., 2019) is shown in magenta. The M-R estimate of the lightest compact object HESS J1731-347  (Horvath et al., 2023) is shown by the shaded area labeled with ‘HESS J1731-347’. The M-R estimates of the two companion stars of the merger event GW170817 are shown by the shaded area labeled with GW170817 M1 (M2).
$m_{s}[{\rm MeV}]$ $a_{4}$ $B_{\rm eff}^{1/4}[{\rm MeV}]$
Set A $150$ $0.8$ $140$
Set B $200$ $0.8$ $140$
Set C $150$ $0.7$ $135$
Set D $200$ $0.7$ $135$
Set E $150$ $0.6$ $130$
Set F $200$ $0.6$ $130$
@@
Table 1: Sets of allowed EoS parameters ($m_{s}$, $c$ and $B_{\rm eff}^{1/4}$) subject to the Bodmer-Witten conjecture.
@@
It is evident that the considered EoS parametrization successfully satisfies key astrophysical observational constraints. The EoS models not only explain the maximum observed pulsar mass but also align well with the M-R posterior distributions of PSR J0740+6620 (Riley et al., 2021) and PSR J0030+0451 (Riley et al., 2019). For the GW170817 merger event, the equivalent M-R posteriors of the component stars, as provided in (Abbott et al., 2018), are derived based on either neutron star EoS models, particularly the piecewise polytropic spectral decomposition approach, or EoS-independent universal relations that were specifically developed for neutron stars. However, in Fig. 2, we present instead the equivalent strange star M-R posteriors, obtained using the mass and tidal deformability measurements for GW170817 from (Abbott et al., 2019) in conjunction with the use of universal relations specifically developed for strange quark stars (Yagi and Yunes, 2017). Additionally, we have included the M-R posterior of the low-mass compact object HESS J1731-347 (Horvath et al., 2023), which was initially proposed as a strong candidate for a strange star.
@@
$\omega/2\pi\,[{\rm kHz}]$ $\omega/2\pi\,[{\rm kHz}]$ $\omega/2\pi\,[{\rm kHz}]$
@@
Set A $2.291$ $2.306$ $2.343$
Set B $2.333$ $2.356$ $2.413$
Set C $2.127$ $2.135$ $2.154$
Set D $2.164$ $2.177$ $2.205$
Set E $1.970$ $1.974$ $1.985$
Set F $2.003$ $2.010$ $2.026$
[…]
II.3 Bulk viscosity in strange quark stars

### v2

@@ I Introduction
All these previous works considered only the hadronic sources of shear viscosity originating from $ee$ scattering as the dominant source of shear viscous damping at low temperatures ($T\ll 10^{10}$ K which is $\approx 1$ MeV) during the inspiral. But, as explained earlier, the densities at the core of neutron stars can be high enough for strange matter such as hyperons or deconfined quark matter to appear as stable components. Recently, Ghosh et al. Ghosh et al. (2024) demonstrated that the bulk viscosity originating from non-leptonic weak processes involving hyperons can be several orders higher ($\approx 10^{8}-10^{10}$ times) than shear viscosity from $ee$ scattering in the temperature range of $10^{6}-10^{8}$ K Lindblom and Owen (2002); Chatterjee and Bandyopadhyay (2006); Alford and Haber (2021), and the dissipation of the dominant $f$-mode introduces an additional phase of the order $10^{-3}-0.5$ rad to the GW signal depending on the neutron star masses, which should in principle be detectable by the next generation GW detectors (Ghosh et al., 2025).
@@
In general, viscous dissipation in strange quark stars is driven by shear and bulk viscosities. Shear viscosity generated by quark-quark scattering due to gluon exchange has been calculated up to $\mathcal{O}(\alpha_{s})$ in a standard Fermi liquid approximation Haensel and Jerzak (1989) and corrected in Heiselberg and Pethick (1993), taking into account that the dynamical screened transverse gluon interactions lead to a non-Fermi liquid behavior. At densities and mode frequencies found in compact stars and at temperatures ranging from $10^{7}$ to $10^{10}$ K, shear viscosity results to be orders of magnitude smaller than the bulk viscosity, the latter being more relevant in the tidal energy dissipation.
@@
This article is structured as follows: in Sec. II we briefly recapitulate the Newtonian tidal heating formalism during the inspiral stage of the coalescence of BNS and estimate the bulk viscous dissipated energy. We also present the non-ideal bag model we use to describe unpaired strange quark matter in subsection III.1, the strange star EoS in subsection III.2, and calculate the bulk viscosity associated with non-leptonic weak interactions among quarks in subsection III.3, which are used to calculate the tidal dissipated energy in the binary inspiral. In Sec. IV, we estimate the temperature reached during the inspiral due to this dissipation from strange quark bulk viscosity. Then in Sec.V, we estimate the additional phase correction introduced in the frequency domain gravitational waveform due to this viscous dissipation and compare with the current limits with future generation GW detector sensitivity to argue about the detectability. In Sec. VI we discuss the main implications of this work and also future directions. We use natural units, assuming $G=\hbar=c=k_{B}=1$, unless explicitly stated otherwise.
@@ II Tidal dissipated energy from mode damping
$\gamma_{\alpha}=\dot{E}_{\text{visc},\alpha}/(2E_{\alpha}),$ (1)
@@
where $\xi_{\alpha}(r)$ is the eigenfunction with the orthogonality and normalization condition
$\int d^{3}x\,\rho\,\xi^{*}_{\alpha}\xi_{\alpha^{\prime}}=\delta_{\alpha\alpha^{\prime}}$ (5)
where $\rho$ is the energy density, and $a_{\alpha}(t)$ is the amplitude of the particular eigenmode due to the tidal field of the companion and as stated before, $\alpha$ is the normal mode index. For a particular mode, the eigenfunction can be written as the sum of the radial $\xi^{r}_{nl}$ and tangential $\xi^{\perp}_{nl}$ components
@@
$\gamma=\frac{1}{2}\frac{(l+|m|)!}{(l-|m|)!}\int_{0}^{R}r^{2}dr\zeta\left(\frac{\partial\xi^{r}}{\partial r}+\frac{2}{r}\xi^{r}-l(l+1)\frac{\xi^{\perp}}{r}\right)^{2},$ (8)
where $\xi^{r}$ and $\xi^{\perp}$ are the radial and perpendicular components of the $f$-mode eigenfunction, and the corresponding velocity field can be written as $\bm{V}=-i\omega\bm{\xi}$, where $\omega$ is the $f$-mode angular frequency, which is determined via a relativistic Cowling approximation Pradhan and Chatterjee (2021) as outlined in appendix A. The normalised mode eigenfunctions are also used to calculate the tidal coupling strength $Q_{nl}$ defined as
$Q_{nl}=\int_{0}^{R}\rho lr^{l+1}[\xi_{nl}^{r}(r)+(l+1)\xi_{nl}^{\perp}(r)]dr,$ (9)
where $R$ is the radius of the star. In the adiabatic limit away from the merger when the orbital evolution is sufficiently slow ($\dot{\Omega}/\Omega\ll 1$ where $\Omega$ is the orbital angular velocity), and far away from the merger and the resonant $f$-mode frequency, we can estimate the energy dissipated in the star with mass $M_{1}$ as Ghosh et al. (2024)
$\dot{E}_{\rm visc}=\frac{24\pi}{5}q^{2}(1+q)\frac{M_{1}^{2}}{R_{1}}\frac{Q_{1}^{2}}{\bar{\omega}_{1}^{4}}\left(\frac{R_{1}}{D}\right)^{9}\gamma,$ (10)
where $M_{1}$ and $qM_{1}$ are the mass of the heavier star and its companion respectively with $q$ being the mass ratio ($q\leq 1$), $Q_{1}$ is the tidal coupling of the $l=m=2$ ff-mode, $R_{1}$ is the radius of the star 1, $\bar{\omega}_{1}$ the normalised (by $\sqrt{M_{1}/R_{1}^{3}}$) angular frequency of the $f$-mode and $D$ is the orbital separation.
III Microphysics of strange stars
III.1 EoS for strange quark matter
We briefly describe the non-ideal bag model that we use to study unpaired strange quark matter ($u,d,s$) with electrons ($e$). The thermodynamic potential $\Omega_{\rm ni}$ (normalised by the volume) as a function of the particles’ chemical potential $\mu_{i}$ with $i=u,d,s,e$ and temperature $T$ can be written as follows
$\Omega_{\rm ni}=\Omega_{e}^{(0)}+a_{4}\sum_{f=u,d,s}\Omega_{f}^{(0)}+B_{\rm eff},$ (11)
@@ III.1 EoS for strange quark matter
where $E_{i,k}=\sqrt{k^{2}+m_{i}^{2}}$, $g_{i}$ is the degeneracy factor and $m_{i}$ is the mass of the particles. For electrons, $g_{e}=2$, considers their spin and for quarks, $g_{f}=2N_{c}$, includes the spin and color charge degrees of freedom being $N_{c}=3$ the number of colors. The mass of the electron is around $0.511$ MeV, but here we consider it to be massless. In this approximation and following the procedure in Ref. Laine and Vuorinen (2016), we get
@@
[…]
@@
and the beta equilibrium conditions
@@
Figure 1: Stability window for the non-ideal bag model. The strange quark mass $m_{s}$ is plotted as a function of $B_{\rm eff}^{1/4}$ for three different values of the nonperturbative parameter $a_{4}$. The shadow regions represent stable strange quark matter according to the Bodmer-Witten conjecture. We have imposed $m_{s}=95$ MeV as the lowest value of the strange quark mass Navas et al. (2024).
III.2 Strange Star EoS and Stellar Properties
@@
In Fig. 1, we plot the regions which satisfy the constraints imposed by the Bodmer-Witten conjecture in the parameter space of $B_{\rm eff}^{1/4}$ and $m_{s}$ as reported in Ref. Torres and Menezes (2013). The vertical lines represent the region where the energy density per baryon for two-flavour quark matter equals $930$ MeV and the other curved lines depict the values where the energy density per baryon for strange quark matter is $930$ MeV. The areas to the left of the vertical lines represent the regions where two-flavour quark matter is stable for a given value of $a_{4}$. The shaded regions show the stability windows setting $m_{u}=m_{d}=4$ MeV and $m_{s}=95$ MeV as the lowest value for the strange-quark mass as many results predicts approximately this value at a renormalization scale of $2$ GeV Navas et al. (2024) and given that this mass is expected to be higher at densities of compact stars as we will discuss below.
In Ref. Fraga et al. (2001) it was found that the perturbative QCD pressure up to $\mathcal{O}(\alpha_{s}^{2})$ for three massless quarks can be approximated by a similar phenomenological bag model like the one in Eq. (11) at chemical potentials from $300$ up to $650$ MeV relevant for quark stars by setting $a_{4}\approx 0.63$ while the bag constant varies according to the renormalization scale choice. In this work, we set $a_{4}=0.8,\,0.7,\,0.6$, which induce small to medium deviations of an ideal free Fermi gas ($a_{4}=1$) and are close to the value reported in Ref. Fraga et al. (2001). For the non-ideal bag model, we find that $m_{s,\text{max}}\approx 310$ MeV is the maximum value allowed for the strange quark mass consistent with the constraints mentioned above.
@@
Figure 2: M-R sequences for the EoS parametrization tabulated in table 1. Horizontal bands correspond to masses $M=2.08{\pm 0.07}M_{\odot}$ of PSR J0740++6620 (Fonseca et al., 2021) and $M=2.01^{+0.04}_{-0.04}M_{\odot}$ of PSR J0348++0432 (Antoniadis et al., 2013). The 90% contour of M-R measurement for PSR J0740+6620 corresponding to Riley et al. (Riley et al., 2021; Salmi et al., 2024) is shown in black, and for PSR J0030+0451 (Riley et al., 2019; Vinciguerra et al., 2024) is shown in magenta. The M-R estimate of the lightest compact object HESS J1731-347 (Doroshenko et al., 2022) is shown by the shaded area labelled with ‘HESS J1731-347’. Although recent works (Alford and Halpern, 2023) puts the original spectral analysis into question. The M-R estimates of the two companion stars of the merger event GW170817 are shown by the shaded area labelled with GW170817 M1 (M2).
According to the stability windows found in Fig. 1, we define the sets of parameters ($m_{s}$, $a_{4}$ and $B_{\rm eff}^{1/4}$) in table 1 so as to determine the EoSs, while assuming $m_{u}=m_{d}=4$ MeV and massless electrons. A variation of the bag constant in the allowed domain does not have a strong impact on the quantities we are interested. In particular the bulk viscosity does not depend on this value. Regarding the strange quark mass, we choose $m_{s}=150,\,200$ MeV given that it is expected to be higher than the values reported in Ref. Navas et al. (2024) at densities a few times the nuclear saturation density, due to the presence of quark antiquark condensates, as it occurs in Nambu-Jona-Lasinio models BUBALLA (2005). Then, we solve the Tolman-Oppenheimer-Volkoff equations (TOV equations) to obtain the mass (M)- radius (R) relations and the $f$-mode oscillation frequencies within the Cowling approximation as described in the appendix A. In Fig. 2, we display the M-R relations obtained with the set of parameters described above. Comparing the curves obtained using sets A-B, C-D, and E-F, where $a_{4}$ and the bag constant are fixed, we infer that decreasing the value of the strange quark mass produces higher maximum masses and radii. Using the EoS of the non-ideal bag model it is possible to produce maximum masses above $2M_{\odot}$ unlike the original MIT bag model, which generates at most a maximum mass up to $1.85M_{\odot}$ Lopes et al. (2021). A similar effect that is more remarkable occurs when we decrease $a_{4}$ Lopes et al. (2021) (see, for example, set A and C, the small variation of the bag constant does not have a significant effect).
Set $m_{s}[{\rm MeV}]$ $a_{4}$ $B_{\rm eff}^{1/4}[{\rm MeV}]$
A $150$ $0.8$ $140$
B $200$ $0.8$ $140$
C $150$ $0.7$ $135$
D $200$ $0.7$ $135$
E $150$ $0.6$ $130$
F $200$ $0.6$ $130$
@@
Table 1: Sets of allowed EoS parameters ($m_{s}$, $a_{4}$ and $B_{\rm eff}^{1/4}$) subject to the Bodmer-Witten conjecture.
@@
It is evident that the considered EoS parametrization successfully satisfies key astrophysical observational constraints. The EoS models not only explain the maximum observed pulsar mass but also align well with the M-R posterior distributions of PSR J0740+6620 (Riley et al., 2021; Salmi et al., 2024) and PSR J0030+0451 (Riley et al., 2019; Vinciguerra et al., 2024). For the GW170817 merger event, the equivalent M-R posteriors of the component stars, as provided in (Abbott et al., 2018), are derived based on either neutron star EoS models, particularly the piecewise polytropic spectral decomposition approach, or EoS-independent universal relations that were specifically developed for neutron stars. However, in Fig. 2, we present instead the equivalent strange star M-R posteriors, obtained using the mass and tidal deformability measurements for GW170817 from (Abbott et al., 2019) in conjunction with the use of universal relations specifically developed for strange quark stars (Yagi and Yunes, 2017). Additionally, we have included the M-R posterior of the low-mass compact object HESS J1731-347 (Horvath et al., 2023), which was initially proposed as a strong candidate for a strange star.
@@
Set $\omega/2\pi\,[{\rm kHz}]$ $\omega/2\pi\,[{\rm kHz}]$ $\omega/2\pi\,[{\rm kHz}]$
@@
A $2.291$ $2.306$ $2.343$
B $2.333$ $2.356$ $2.413$
C $2.127$ $2.135$ $2.154$
D $2.164$ $2.177$ $2.205$
E $1.970$ $1.974$ $1.985$
F $2.003$ $2.010$ $2.026$
Table 2: $f$-mode frequencies $\omega/2\pi$ in kHz at $M/M_{\odot}=1.4,\,1.6,\,1.8$ for the sets defined in table 1.
III.3 Bulk viscosity in strange quark stars